OpenStudy (anonymous):

what is the square root of 216k^2

OpenStudy (anonymous):

square root of 216 is 8 sum yea cause ive been trin to work it out uhhhh!...

OpenStudy (anonymous):

\[\sqrt{216k^{2}} = \sqrt{36*6*k^{2}}=\sqrt{6^{2}*6*k^{2}}\]

OpenStudy (anonymous):

\[6k \sqrt{6}\]